# Flock: A Knowledge Graph Foundation Model via Learning on Random Walks

### v1

@@ Abstract
We study the problem of zero-shot link prediction on knowledge graphs (KGs), which requires models to generalize over novel entities and novel relations. Knowledge graph foundation models (KGFMs) address this task by enforcing equivariance over both nodes and relations, learning from structural properties of nodes and relations, which are then transferable to novel graphs with similar structural properties. However, the conventional notion of deterministic equivariance imposes inherent limits on the expressive power of KGFMs, preventing them from distinguishing structurally similar but semantically distinct relations. To overcome this limitation, we introduce probabilistic node-relation equivariance, which preserves equivariance in distribution while incorporating a principled randomization to break symmetries during inference. Building on this principle, we present Flock, a KGFM that iteratively samples random walks, encodes them into sequences via a recording protocol, embeds them with a sequence model, and aggregates representations of nodes and relations via learned pooling. Crucially, Flock respects probabilistic node-relation equivariance and is a universal approximator for isomorphism-invariant link-level functions over KGs. Empirically, Flock perfectly solves our new diagnostic dataset Petals where current KGFMs fail, and achieves state-of-the-art performances on entity- and relation prediction tasks on 54 KGs from diverse domains.
@@ 1 Introduction
Knowledge graph foundation models (KGFMs) (Lee et al., 2023; Geng et al., 2023; Galkin et al., 2024; Zhang et al., 2024; Cui et al., 2024; Huang et al., 2025) aim to infer missing links over novel knowledge graphs (KGs) that are not part of the training graphs or domains. This task requires generalization to both unseen nodes and unseen relation types. To achieve this, KGFMs rely on learning node and relation invariants: structural properties of nodes and relations that are transferable across KGs even when their relational vocabularies differ. This inductive bias is formalized by Gao et al. (2023) as double-equivariance — equivariance under permutations of both entities and relations — and used as a core principle in the design of KGFMs in the literature.
@@
Figure 1: A KG representing characters’ relationships in Star Wars movies. Blue arrows indicate $\mathsf{like}$, red arrows – $\mathsf{dislike}$, and green arrows indicate relation ($\mathsf{friendWith}$).
@@
Problem statement. In this work, we challenge the fundamental assumption of existing KGFMs dictated by strict equivariance: structural isomorphism of relations implies semantic equivalence. Consider, for example the KG from Figure 1, where the relations like and dislike are structurally isomorphic, and yet they represent semantically opposite relations. In this motivating example, any KGFM that computes relation invariants is forced to assign the same representation to both like and dislike — losing the ability to distinguish between two entities with opposite relationships. This expressiveness limitation is an architectural one and cannot be resolved through finetuning, which further limits the downstream use of existing KGFMs. This raises a central question: How to design KGFMs that are both expressive and have the right inductive bias for generalization?
@@
Our approach. We propose a new approach for KGFMs, which relies on probabilistic node-relation equivariance as inductive bias. Instead of enforcing deterministic equivariance over nodes and relations, these KGFMs respect probabilistic node-relation equivariance. This relaxes the hard constraint that “structurally isomorphic relations must have identical representations”, and requires only that “the representations of structurally isomorphic relations need to be equivalent in distribution” over a model’s stochastic processes. This way, the model retains crucial inductive bias needed for generalizing across different KGs, while the stochasticity of each forward pass ensures that structurally identical but semantically distinct relations are assigned different representations, allowing the model to distinguish between them (Srinivasan & Ribeiro, 2020; Abboud et al., 2021).
@@
Inspired by the success of models that learn probabilistic invariants via random walks (Perozzi et al., 2014; Grover & Leskovec, 2016; Nikolentzos & Vazirgiannis, 2020; Kim et al., 2025), we introduce Flock, a KGFM that inherently computes probabilistic node-relation invariants. Given a (potentially unseen) KG, and a query, in each iteration, Flock first samples a set of random walks over KG based on the query, noting down both encountered nodes and relations with a recording protocol. To ensure the model can generalize to unseen entities and relation types, we anonymize all nodes and relations, enforcing that Flock only learn from their structural roles. These anonymized sequences are then fed into a sequence processor, and the representations for each node and relation are aggregated via a consensus protocol. Finally, we construct per-query (triple) features from the aggregated entity and relation embeddings and input them into a binary classifier for link prediction.
@@
Key findings and contributions. The design of Flock offers several key advantages over existing KGFMs. First, it entirely abandons the conventional two-stage process of encoding relations and node representations via two separate networks, and does not rely on message-passing at all, thereby avoiding the well-known expressivity limitations of MPNNs on KGs (Barceló et al., 2022; Huang et al., 2023; 2025). Second, Flock is a universal approximator (see Proposition 4.1), and thus can approximate every link-level function defined on KGs of any bounded size. Finally, Flock architecture inherently respects the principle of probabilistic node-relation equivariance, which enables a strong generalization capacity. Our experimental results over both entity prediction and relation prediction validate the strength of this approach, demonstrating that Flock consistently outperforms state-of-the-art KGFMs on existing benchmarks. Our contributions can be summarized as follows:
@@
  • We highlight a limitation in existing KGFMs: their over-reliance on deterministic node–relation equivariance prevents them from distinguishing between structurally similar but semantically different relations, limiting their expressivity.
@@
  • We introduce probabilistic node-relation equivariance, a property for KGFMs that ensures invariance only in distribution, thus balancing the model expressivity and generalization.
@@
  • We propose Flock, a KGFM that respects probabilistic node-relation equivariance. Flock replaces the conventional two-stage, message-passing paradigm with a direct sequence encoding approach based on random walks, and acts as a universal approximator of link-level functions.
@@
  • We validate our approach on both entity and relation prediction tasks across 54 diverse KGs, where Flock consistently achieves state-of-the-art performance over existing KGFMs. We further design a synthetic dataset Petals to confirm our theoretical results empirically.
All proofs of the theoretical results are provided in Appendix C. The code is available at https://github.com/jw9730/flock-pytorch.
@@ 2 Related work
Link prediction and KGFMs. Early methods for inferring missing links in KGs  (Bordes et al., 2013; Sun et al., 2019; Balazevic et al., 2019; Abboud et al., 2020; Schlichtkrull et al., 2018; Vashishth et al., 2020) rely on learned embeddings, hence operating in the transductive setting, incapable of generalizing to unseen entities or relation types. Later GNN-based approaches based on the labeling trick (Teru et al., 2020; Zhang et al., 2021) or conditional message passing (Zhu et al., 2021; 2023; Zhang & Yao, 2022; Zhang et al., 2023b; Huang et al., 2023), unlocked the node inductive scenario, while remaining restricted to a fixed relational vocabulary. KGFMs eliminate this restriction and enable node-relation inductive link prediction over both unseen nodes and relation types through the use of a two-stage process by first encoding relations and then nodes. The first examples of this paradigm are InGram (Lee et al., 2023) and ULTRA (Galkin et al., 2024). Their ideas were extended by TRIX (Zhang et al., 2024) to build a more expressive framework. KG-ICL (Cui et al., 2024) achieved full inductivity by combining in-context learning with node-relation tokenization. ISDEA (Gao et al., 2023) and MTDEA (Zhou et al., 2023) highlighted the benefits of equivariance over both nodes and relations. MOTIF (Huang et al., 2025) was proposed as a general KGFM framework, supported by a theoretical analysis of the expressive power of KGFMs. Our work further advances the field with the first stochastic KGFM, which is invariant in probability and provably more expressive than all the existing methods. Notably, Flock achieves universality without any form of message passing, instead relying on random walks and sequence models to encode both nodes and relations anonymously to ensure generalization.
@@
Random walks for graph representations. Random walks have attracted a lot of attention in graph learning, due to their simplicity and ability to gather context from neighborhoods. DeepWalk (Perozzi et al., 2014) and node2vec (Grover & Leskovec, 2016) were among the first to explore the potential of random walks for producing graph embeddings, treating walks as analogues of sentences in natural language and processing them using skip-gram models. Nikolentzos & Vazirgiannis (2020) generated graph-level task predictions by executing joint random walks on direct products of graphs with their extracted subgraphs. CRaWL (Tönshoff et al., 2021) represents the input graph as a collection of random walks and processes them with a 1-dimensional convolutional NN. WalkLM (Tan et al., 2023) samples random walks from graphs with textual features, passing them to a language model for embedding generation. RWNN (Kim et al., 2025) and RUM (Wang & Cho, 2024) anonymize the extracted walks and process them with sequence models and RNNs, respectively. NeuralWalker (Chen et al., 2025) aggregates embeddings derived by encoding random walks into message passing layers.
@@
Probabilistic invariance. Neural architectures that enforce invariance to specific transformations often exhibit more stable training and improved performance (Bronstein et al., 2021), but this inductive bias can reduce their expressivity by preventing the model from distinguishing non-equivalent inputs. In graph learning, this trade-off is exemplified by MPNNs, whose power is limited by the 1-WL test (Xu et al., 2019; Morris et al., 2019). Randomization has emerged as a solution, enhancing expressivity through techniques such as noise injection (Abboud et al., 2021), vertex dropping (Papp et al., 2021), subgraph sampling (Bevilacqua et al., 2022; Zhang et al., 2023a), dynamic rewiring (Finkelshtein et al., 2024), and random walks (Kim et al., 2025; Wang & Cho, 2024). Despite their stochasticity, such methods can remain probabilistically invariant, ensuring that equivalent inputs yield identical expected outputs, or even identical output distributions. We extend the notion of probabilistic invariance to KGs and prove that Flock satisfies invariance in distribution.
@@ 3 Preliminary
Knowledge graphs. A knowledge graph (KG) is a tuple $G=(V,E,R)$, where $V$ denotes the set of entities (nodes), $R$ the set of relation types, and $E\subseteq V\times R\times V$ the set of labeled edges (facts). A fact is written as $(h,r,t)$ (or $h\xrightarrow{r}t$ interchangeably) with $r\in R$ and $h,t\in V$. A (potential) link in $G$ is any triple $(h,r,t)$ in $V\times R\times V$, regardless of whether it is present in $E$. We denote by $R^{-1}$ the set of inverses of relations $R$, defined as $\{r^{-1}\mid r\in R\}$, and mean $r$ when writing $(r^{-1})^{-1}$.
@@
Probabilistic invariance. Let $\mathbb{K}_{n,m}$ be the space of knowledge graphs with $n$ vertices and $m$ relation types. A stochastic KG model $\varphi$ can be viewed as a function that takes a KG $G$ as the input and returns a random variable $\varphi(G)$. Following Kim et al. (2025), we call $\varphi$ invariant in probability if
$\forall G,G^{\prime}\in\mathbb{K}_{n,m}:\qquad G\simeq G^{\prime}\implies\varphi(G)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\varphi(G^{\prime})$
i.e. the distributions of $\varphi(G)$ and $\varphi(G^{\prime})$ are equal. In particular, this implies $\mathbb{E}[\varphi(G)]=\mathbb{E}[\varphi(G^{\prime})]$.
@@ 4 Methodology
Figure 2: Overall pipeline of Flock. In each updating step, Flock samples random walks on the KG, anonymizes the encountered nodes and relations via a recording protocol, and feeds the sequences in a sequence processor to compute node and relation representations. A consensus protocol then pools them back to the original KG’s nodes and relations.
We present Flock, a KGFM respecting probabilistic node-relation invariance. Flock is a randomized function $X_{\theta}(\cdot)$ which takes as input a KG $G=(V,E,R)$ and a link prediction query $q$. We consider two types of queries: entity prediction $q=(h,r,?)$ and relation prediction $q=(h,?,t)$. Flock outputs a random variable $\hat{{\bf y}}\sim X_{\theta}(G,q)$ which is suited for the task at hand. For entity prediction, it outputs $\hat{{\bf y}}:V\to[0,1]$ such that a potential link $(h,r,t)$ can be evaluated by $\hat{{\bf y}}(t)\in[0,1]$. For relation prediction, it outputs $\hat{{\bf y}}:R\to[0,1]$ such that a link $(h,r,t)$ can be evaluated by $\hat{{\bf y}}(r)$. In practice, we often ensemble repeated predictions $\hat{\bf y}_{1},...,\hat{\bf y}_{P}\sim X_{\theta}(G,q),P\in\mathbb{N}$ via averaging.
We describe the architecture of Flock in Section 4.1 focused on four main components, and then analyze its theoretical properties in Section 4.2, showing universality and probabilistic equivariance.
@@ 4.1 Flock
${\bf v}^{(i+1)}\coloneqq{\bf v}^{(i)}+\Delta{\bf v},\quad{\bf r}^{(i+1)}\coloneqq{\bf r}^{(i)}+\Delta{\bf r},\quad(\Delta{\bf v},\Delta{\bf r})\sim{\rm update}_{\theta_{i}}({\bf v}^{(i)},{\bf r}^{(i)}).$ (1)
@@
    Random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $l$ on the input KG.
@@
Random walks. In Flock, random walks are central in two ways: they rewrite the connectivity of nodes and relations as sequences, and support generalization via probabilistic equivariance.
Formally, the random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $l$ on KG $G$. Each random walk $\eta$ is a chain of random variables, written as:n as:
$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{l}}v_{l},\quad v_{s}\in V,r_{s}\in R,(v_{s-1},r_{s},v_{s})\in E,$ (2)
where the underlying transition mechanism and $l$ are hyperparameters.
@@
$\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\cdots\xrightarrow{\phi(r_{l})}\pi(v_{l})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\cdots\xrightarrow{s_{l}}u_{l}$ (3)
where $v_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{l}}v_{l}$ and $u_{0}\xrightarrow{s_{1}}\cdots\xrightarrow{s_{l}}u_{l}$ follow the distributions of $\eta(G,l)$ and $\eta(H,l)$, respectively. In such case, the isomorphism $(\pi,\phi)$ yields a natural translation from walks in $G$ to $H$.
In Flock, we use a simple random walk algorithm, which we show to be invariant in probability, while working robustly in practice. Specifically, given a base walk count $n$, for entity prediction queries $(h,r,?)$, we use $3n$ walks and choose three types of start locations. We start $n$ walks at query node $h$ by fixing $v_{0}=h$, start $n$ walks by choosing a random relation $s$, then a random edge $(v,s,u)\in E$, and then fixing the first step as $v\xrightarrow{s}u$, and start the rest of $n$ walks at random nodes in $V$. For relation prediction queries $(h,?,t)$, we additionally start $n$ walks at the tail node $t$ by fixing $v_{0}=t$, sampling a total of $4n$ walks.
For the transition mechanism, we use uniform walks with non-backtracking, with minor modifications to handle directed multi-edges of KGs. Despite the simplicity, we find that this choice works well in practice, consistent with findings of prior work (Tönshoff et al., 2021; Kim et al., 2025).
@@
$n=n_{\rm train}\times{\rm harmonic}\ {\rm mean}\left(\frac{|V|}{|V|_{\rm train}},\frac{|E|}{|E|_{\rm train}}\right)$ (4)
where $|V|_{\rm train},|E|_{\rm train}$ are average numbers of nodes and edges in pretraining KGs, respectively. Intuitively, this scales $n$ proportionally to the sizes of test KGs relative to pretraining. In practice, we clamp $n$ to the nearest power of $2$ in an interval to keep GPU memory usage in a range.
@@
$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}v_{2}\xrightarrow{r_{1}^{-1}}v_{0}\qquad\mapsto\qquad 1\xrightarrow{\alpha}2\xrightarrow{\beta}3\xrightarrow{\alpha^{-1}}1,$ (5)
@@
Now that the recordings $\bf z$} only encode structural information of KG, we can safely process them with an arbitrary neural network $f_{\theta}:{\bf z}\mapsto{\bf h}$ without the risk of losing invariance. Since $\bf z$} are sequences, we choose sequence networks to leverage their inductive bias. Specifically, we use bidirectional GRU (Cho et al., 2014) equipped with RMSNorm (Zhang & Sennrich, 2019) and SwiGLU feedforward network (Shazeer, 2020), which provided robust results.
@@
$f_{\theta}:{\bf z}\mapsto{\bf h}=(\Delta{\bf v}_{0},a_{0})\xrightarrow{\Delta{\bf r}_{1},b_{1}}(\Delta{\bf v}_{1},a_{1})\xrightarrow{\Delta{\bf r}_{2},b_{2}}\cdots.$ (7)
@@
After sequence processing, we are left with a handful of state update proposals ${\bf h}_{1:N}$ from $f_{\theta}$, that are positionally aligned with random walks $\eta_{1:N}$ on KG $G=(V,E,R)$. The consensus protocol $c$ uses the information to decide final state updates $\Delta{\bf v}:V\to\mathbb{R}^{d}$ and $\Delta{\bf r}:R\to\mathbb{R}^{d}$.

### v2

@@ Abstract
We study the problem of zero-shot link prediction on knowledge graphs (KGs), which requires models to generalize to novel entities and novel relations. Knowledge graph foundation models (KGFMs) address this task by enforcing equivariance over both nodes and relations, which enables them to learn structural properties of nodes and relations that transfer to novel KGs with similar structure. However, the conventional notion of deterministic equivariance inherently limits the expressive power of KGFMs, as it prevents them from distinguishing relations that are structurally similar but semantically distinct. To overcome this limitation, we propose to leverage probabilistic node-relation equivariance, which preserves equivariance in distribution while using structured randomness to break symmetries at inference time. Building on this principle, we present Flock, a KGFM that iteratively samples random walks, encodes them into sequences, embeds them with a sequence model, and aggregates node and relation representations through learned pooling. Flock respects probabilistic node-relation equivariance and, crucially, is a universal approximator for isomorphism-invariant link-level functions over KGs. Empirically, Flock perfectly solves our new diagnostic dataset Petals on which current KGFMs fail, and achieves state-of-the-art performance on entity and relation prediction tasks across 54 KGs from diverse domains.
@@ 1 Introduction
Knowledge graph foundation models (KGFMs) (Galkin et al., 2024; Zhang et al., 2024; Huang et al., 2025) aim to infer missing links over novel knowledge graphs (KGs) that are not part of the training data or domains. This task requires generalization to both unseen nodes and unseen relation types. To achieve this, KGFMs learn node and relation invariants: structural properties of nodes and relations that are transferable across KGs even when their relational vocabularies differ. This inductive bias is formalized as double equivariance (Gao et al., 2023)—equivariance under permutations of both entities and relations—and used as a core design principle of current KGFMs.
@@
Figure 1: A KG representing characters’ relationships in Star Wars movies. Blue arrows indicate $\mathsf{like}$, red – $\mathsf{dislike}$, and green arrows indicate $\mathsf{friendWith}$.
Problem statement.
@@
In this work, we challenge a fundamental assumption dictated by strict equivariance in existing KGFMs: structural isomorphism of relations implies semantic equivalence. Consider, for example, the KG in Figure˜1, where the relations $\mathsf{like}$ and $\mathsf{dislike}$ are structurally isomorphic yet semantically opposite. Any KGFM that computes relation invariants is forced to assign the same representation to both $\mathsf{like}$ and $\mathsf{dislike}$—losing the ability to distinguish entities with opposite relationships. This expressiveness limitation is an architectural one and cannot be resolved through finetuning, further limiting the downstream use of existing KGFMs. This raises a central question: how can we design KGFMs that are both expressive and have the right inductive bias for generalization?
Approach.
@@
We propose a new approach for KGFMs that relies on probabilistic node-relation equivariance as inductive bias, instead of enforcing deterministic equivariance over nodes and relations. Probabilistic equivariance relaxes the hard constraint that “structurally isomorphic relations must have identical representations”, and requires only that the representations of structurally isomorphic relations need to be equivalent in distribution over a model’s stochastic predictive processes (Srinivasan & Ribeiro, 2020; Abboud et al., 2021). This way, the model retains the inductive bias needed for generalizing across different KGs, while the stochasticity of each forward pass ensures that structurally identical but semantically distinct relations are assigned different representations.
@@
Inspired by the success of models that learn probabilistic invariants via random walks (Perozzi et al., 2014; Grover & Leskovec, 2016; Nikolentzos & Vazirgiannis, 2020; Kim et al., 2025), we introduce Flock, a KGFM that inherently computes probabilistic node-relation invariants. Given a potentially unseen KG and a query, Flock iteratively samples random walks over the KG based on the query, encoding both encountered nodes and relations with a recording protocol. To ensure the model can generalize to unseen entities and relation types, we anonymize all nodes and relations, enforcing that Flock only learn from their structural roles. These anonymized sequences are then fed into a sequence processor, and the representations for each node and relation are aggregated via a consensus protocol. Finally, we construct per-query features from the aggregated entity and relation embeddings and input them into a binary classifier for link prediction.
Key findings and contributions.
@@
The design of Flock offers several key advantages over existing KGFMs. First, it entirely abandons the conventional two-stage process of encoding relations and node representations via two separate networks, and does not rely on message passing at all, thereby avoiding the well-known expressivity limitations of MPNNs on KGs (Barceló et al., 2022; Huang et al., 2023; 2025). Second, Flock is a universal approximator (Proposition 4.1), capable of approximating every link-level function on KGs of any bounded size. Finally, Flock’s architecture inherently respects probabilistic node-relation equivariance, enabling strong generalization. Our experiments on both entity and relation prediction validate this approach, demonstrating that Flock consistently outperforms state-of-the-art KGFMs on existing benchmarks. Our contributions are:
@@
  • We highlight a limitation of existing KGFMs: their reliance on deterministic node–relation equivariance prevents them from distinguishing between structurally similar but semantically different relations, limiting their expressivity.
@@
  • We propose to leverage probabilistic node-relation equivariance, a property for KGFMs that ensures equivariance only in distribution, as an effective solution balancing the model expressivity and generalization.
@@
  • We propose Flock, a KGFM that respects probabilistic node-relation equivariance. Flock replaces the conventional two-stage, message-passing paradigm with a direct sequence encoding approach based on random walks, and is a universal approximator of link-level functions.
@@
  • We validate our approach on entity and relation prediction tasks across 54 diverse KGs, where Flock consistently achieves state-of-the-art performance. We further construct a synthetic dataset Petals to confirm our theoretical results empirically.
All proofs are in Appendix˜B. The code is available at https://github.com/jw9730/flock.
@@ 2 Related work
Early methods for inferring missing links in KGs relied on learned embeddings (Bordes et al., 2013; Sun et al., 2019; Balazevic et al., 2019; Abboud et al., 2020; Schlichtkrull et al., 2018; Vashishth et al., 2020), operating in the transductive setting and incapable of generalizing to unseen entities or relation types. Later MPNN-based methods based on the labeling trick (Teru et al., 2020; Zhang et al., 2021) or conditional message passing (Zhu et al., 2021; 2023; Zhang & Yao, 2022; Zhang et al., 2023b; Huang et al., 2023) unlocked the node-inductive scenario, while remaining restricted to a fixed relational vocabulary. KGFMs eliminate this restriction and enable node-relation inductive link prediction over both unseen nodes and relations (Geng et al., 2023), typically by first encoding relations and then nodes. Early examples are InGram (Lee et al., 2023) and ULTRA (Galkin et al., 2024), later extended by TRIX (Zhang et al., 2024) into a more expressive framework. KG-ICL (Cui et al., 2024) achieved full inductiveness by combining in-context learning with node-relation tokenization. ISDEA (Gao et al., 2023) and MTDEA (Zhou et al., 2023) highlighted the benefits of node-relation equivariance, while MOTIF (Huang et al., 2025) proposed a general KGFM framework with a theoretical analysis of their expressive power. Our work advances the field with a stochastic KGFM that achieves invariance in probability to ensure generalization while being provably more expressive than the existing methods. Notably, Flock achieves universality without any form of message passing, instead relying on random walks and sequence models to encode both nodes and relations anonymously. This is distinct from prior stochastic KGFMs that rely on random initialization of message passing (Lee et al., 2023; Gao et al., 2023).
Random walks for graph representations.
@@
Random walks have been widely adopted in graph learning due to their simplicity and ability to gather context from neighborhoods. DeepWalk (Perozzi et al., 2014) and node2vec (Grover & Leskovec, 2016) were among the first approaches, treating walks as analogues of natural language sentences and processing them with skip-gram models. Subsequent work has developed diverse neural architectures based on random walks: Nikolentzos & Vazirgiannis (2020) generates graph-level predictions using joint walks on direct products of graphs and their subgraphs, CRaWL (Tönshoff et al., 2021) represents a graph as a collection of walks and processes them with a convolutional network, WalkLM (Tan et al., 2023) samples walks from graphs with textual features and embeds them using a language model, RWNN (Kim et al., 2025) and RUM (Wang & Cho, 2024) anonymize walks and process them with sequence networks, and NeuralWalker (Chen et al., 2025) integrates walk encodings into message passing.
Probabilistic invariance.
@@
Neural architectures that enforce invariance to specific transformations often exhibit more stable training and improved performance (Bronstein et al., 2021), but this inductive bias can reduce their expressivity by preventing the model from distinguishing non-equivalent inputs. In graph learning, this trade-off is exemplified by MPNNs, whose power is limited by the 1-WL test (Xu et al., 2019; Morris et al., 2019). Randomization has emerged as a solution, enhancing expressivity through techniques such as noise injection (Abboud et al., 2021), vertex dropping (Papp et al., 2021), subgraph sampling (Bevilacqua et al., 2022; Zhang et al., 2023a), dynamic rewiring (Finkelshtein et al., 2024), and random walks (Kim et al., 2025; Wang & Cho, 2024). Despite their stochasticity, such methods can remain probabilistically invariant, ensuring that equivalent inputs yield identical expected outputs, or even identical output distributions. In line with a prior work (Gao et al., 2023), we extend the notion of probabilistic invariance to KGs and prove that Flock satisfies invariance in distribution, and further clarify its usefulness in KG learning.
@@ 3 Preliminary
Knowledge graphs. A knowledge graph (KG) is a tuple $G=(V,E,R)$, where $V$ denotes the set of entities (nodes), $R$ the set of relation types, and $E\subseteq V\times R\times V$ the set of labeled edges (facts). A fact is written as $(h,r,t)$ (or $h\xrightarrow{r}t$ interchangeably) with $r\in R$ and $h,t\in V$. A (potential) link in $G$ is any triple $(h,r,t)$ in $V\times R\times V$, regardless of whether it is present in $E$. We denote by $R^{-1}$ the set of inverses of relations $R$, defined as $\{r^{-1}\mid r\in R\}$, and mean $r$ when writing $(r^{-1})^{-1}$. Further, let $\mathbb{K}_{n,m}$ be the space of knowledge graphs with $n$ vertices and $m$ relation types.
@@
Probabilistic invariance. A stochastic KG model $\varphi$ can be viewed as a function that takes a KG and returns a random variable $\varphi(G)$. Following Kim et al. (2025), we call $\varphi$ invariant in probability if
$\forall G,G^{\prime}\in\mathbb{K}_{n,m}:\qquad G\simeq G^{\prime}\implies\varphi(G)\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\varphi(G^{\prime}),$
i.e., the distributions of $\varphi(G)$ and $\varphi(G^{\prime})$ are equal. In particular, this implies $\mathbb{E}[\varphi(G)]=\mathbb{E}[\varphi(G^{\prime})]$.
@@ 4 Methodology
Figure 2: An overview. In each updating step, Flock (1) samples random walks on the KG (two walks indicated by red and teal, respectively), (2) anonymizes the encountered nodes and relations via a recording protocol (for each walk, nodes are anonymized as $1,2,$... and relations as $\alpha,\beta,$...), and (3) feeds the sequences in a sequence processor to compute node and relation representations. (4) A consensus protocol then pools them back to the original KG’s nodes and relations.
We present Flock, a KGFM respecting probabilistic node-relation invariance. Flock is a randomized function $X_{\theta}(\cdot)$ which takes as input a KG $G=(V,E,R)$ and a link prediction query $q$. We consider two types of queries: entity prediction $q=(h,r,?)$ and relation prediction $q=(h,?,t)$. Flock outputs a random variable $\hat{{\bf y}}\sim X_{\theta}(G,q)$ which is suited for the task at hand. For entity prediction, it outputs $\hat{{\bf y}}:V\to[0,1]$ such that a potential link $(h,r,t)$ can be evaluated by $\hat{{\bf y}}(t)\in[0,1]$. For relation prediction, it outputs $\hat{{\bf y}}:R\to[0,1]$ such that a link $(h,r,t)$ can be evaluated by $\hat{{\bf y}}(r)$.
At test time, we average multiple ($P$) independent stochastic predictions to produce the final output. This improves performance and reduces variance through an ensembling effect.
We describe the architecture of Flock in Section˜4.1 focused on four main components, and then analyze its theoretical properties in Section˜4.2, showing universality and probabilistic equivariance. An expansion on the model details can be found in Appendix˜A.
@@ 4.1 Flock
${\bf v}^{(i+1)}\coloneqq{\bf v}^{(i)}+\Delta{\bf v},\qquad{\bf r}^{(i+1)}\coloneqq{\bf r}^{(i)}+\Delta{\bf r},\qquad(\Delta{\bf v},\Delta{\bf r})\sim{\rm update}_{\theta_{i}}({\bf v}^{(i)},{\bf r}^{(i)}).$
@@
    Random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $\ell$ on the input KG.
@@
Random walks.
In Flock, random walks are central in two ways: they rewrite the connectivity of nodes and relations as sequences, and support generalization via probabilistic equivariance.
Formally, the random walk algorithm produces $n$ random walks $\eta_{1},...,\eta_{n}$ of length $\ell$ on KG $G$. Each random walk $\eta$ is a chain of random variables, written as:n as:
$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{\ell}}v_{\ell},\qquad v_{s}\in V,r_{s}\in R,(v_{s-1},r_{s},v_{s})\in E,$
where the underlying transition mechanism and $\ell$ are hyperparameters.
@@
$\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\cdots\xrightarrow{\phi(r_{\ell})}\pi(v_{\ell})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}u_{0}\xrightarrow{s_{1}}u_{1}\xrightarrow{s_{2}}\cdots\xrightarrow{s_{\ell}}u_{\ell},$
where $v_{0}\xrightarrow{r_{1}}\cdots\xrightarrow{r_{\ell}}v_{\ell}$ and $u_{0}\xrightarrow{s_{1}}\cdots\xrightarrow{s_{\ell}}u_{\ell}$ follow the distributions of $\eta(G,\ell)$ and $\eta(H,\ell)$, respectively. In such case, the isomorphism $(\pi,\phi)$ yields a natural translation from walks in $G$ to $H$.
In Flock, we use a simple random walk algorithm which we show to be invariant in probability. Specifically, we use uniform walks with non-backtracking, with minor modifications to handle directed multi-edges of KGs. Despite the simplicity, we find that this choice works well in practice, consistent with findings of prior works (Tönshoff et al., 2021; Kim et al., 2025).
Under this choice, we diversify the starting locations of walks such that local context around the query $q$ and broad regions of the nodes and relations in a KG are both well-captured. Our diversification strategy is as follows: given a base walk count $n$, for entity prediction queries $(h,r,?)$, we use $3n$ walks with three types of start locations. The first $n$ walks start at query node $h$, capturing local context around the query; the second $n$ walks start by traversing a random edge $(v,s,u)$ where $s$ is a uniformly chosen relation, broadly capturing the relations of the KG including $r$; the last $n$ walks start at random nodes, broadly capturing various regions of the KG. For relation prediction queries $(h,?,t)$, we additionally start $n$ walks at the tail node $t$, sampling a total of $4n$ walks.
@@
$n=n_{\rm train}\times{\rm harmonic}\ {\rm mean}\left(\frac{|V|}{|V|_{\rm train}},\frac{|E|}{|E|_{\rm train}}\right),$ (1)
where $|V|_{\rm train},|E|_{\rm train}$ are average numbers of nodes and edges in pretraining KGs, respectively. Intuitively, this scales $n$ proportionally to the sizes of test KGs relative to pretraining. In practice, we clamp $n$ to the nearest power of $2$ and limit its value in an interval to avoid out-of-memory errors.
@@
$\eta=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}v_{2}\xrightarrow{r_{1}^{-1}}v_{0}\qquad\mapsto\qquad 1\xrightarrow{\alpha}2\xrightarrow{\beta}3\xrightarrow{\alpha^{-1}}1,$
@@
Now that the recordings $\bf z$} only encode structural information of KG, we can safely process them with an arbitrary neural network $f_{\theta}:{\bf z}\mapsto{\bf h}$ without the risk of losing invariance. Since $\bf z$} are sequences, we choose sequence networks to leverage their inductive bias. Specifically, we use bidirectional GRU (Cho et al., 2014) equipped with RMSNorm (Zhang & Sennrich, 2019) and SwiGLU feedforward network (Shazeer, 2020), which provided robust results. To convert anonymizations into input feature vectors to the GRU, we use trainable embedding tables.
@@
$f_{\theta}:{\bf z}\mapsto{\bf h}=(\Delta{\bf v}_{0},a_{0})\xrightarrow{\Delta{\bf r}_{1},b_{1}}(\Delta{\bf v}_{1},a_{1})\xrightarrow{\Delta{\bf r}_{2},b_{2}}\cdots.$
@@
After sequence processing, we are left with a handful of state update proposals ${\bf h}_{1:N}$ from $f_{\theta}$, that are positionally aligned with random walks $\eta_{1:N}$ on KG. The consensus protocol $c$ uses the information to decide final state updates $\Delta{\bf v}:V\to\mathbb{R}^{d}$ and $\Delta{\bf r}:R\to\mathbb{R}^{d}$.